Circular No. 8656

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2005 XA_{54} (LONEOS-HILL) A new comet (with a well-developed tail ≈ 20′′ long to the west-

northwest) discovered by R. Hill on Catalina survey images taken on Jan. 6 was found by T. Spahr, Minor Planet Center, to be identical with an object found by LONEOS on Dec. 4 and designated 2005 XA_{54} (with two nights of observations on MPS 155505). Additional astrometry and orbital elements (T = 2006 Mar. 8.04 TT, q = 1.7764 AU, ω = 15^{o}.45, Ω = 144^{o}.23, i = 16^{o}.90, equinox 2000.0, P = 15.1 years) appear on MPEC 2006-A26.

2005/6 UT  α_{2000}  δ_{2000}  Mag. Observer
Dec. 4.37744 9^{h}20^{m}54.^{s}64 +3^{o}00′13′′.2 18.5 LONEOS
Jan. 6.41154 9 56 51.97 +3 42 03.8 17.1 Hill
(60558) 2000 EC_{98}

Y.-J. Choi and P. R. Weissman, Jet Propulsion Laboratory, report the detection of a cometary coma on B, V, R, and I images of the Centaur object (60558) taken on 2005 Dec. 30.50 UT with the Palomar 5-m reflector (+ Large Format Camera); a 600-second R-band exposure shows a coma of total mag 17.5 extending ∼ 20′′ from the nuclear condensation. Confirming images of the cometary nature were obtained by Choi on 2006 Jan. 1 and by D. Polishook on images obtained at Wise Observatory on Jan. 2. The object was discovered by Spacewatch (discovery observation below from MPEC 2000-E64 and MPS 11487).

2000 UT  α_{2000}  δ_{2000}  Mag.
Mar. 3.41164  11^{h}10^{m}57.^{s}58  +4^{o}40′25′′.0  21.0

SUPERNOVA 2006A IN NGC 7753 Further to IAUC 8646, T. Boles the discovery, on an unfiltered CCD

image taken on Jan. 2.856 UT, of an apparent supernova (mag 18.1) located at α = 23^{h}47^{m}03^{s}.78, δ = +29^{o}28′55′′.7, which is ≈ 13′′.6 west and 5′′.1 south of the center of NGC 7753. W. Li, confirmed 2006A on a Lick Observatory image from Jan. 5.108 (Boles measuring position end figures 03^{s}.81, 55′′.6 from the KAIT image). Traces of 2006A are also visible on an unfiltered CCD image taken by G. Rogers (Long Crendon, England, 0.40-m reflector) on Jan. 4.794. Nothing is visible at this position on Boles’ images from 2005 Sept. 24 and Oct. 10 (limiting mag 19.5) or on Digitized Sky Survey images from 1991 (limiting red mag 21.0) and 1993 (limiting blue mag 20.5).

2006 January 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green